# On the closed neighborhood ideal of the square of the path graph

### v1

@@ Introduction
In 2020, Sharifan and Moradi defined the notion of closed neighborhood ideal of a graph as being the squarefree monomial ideal whose generators are given by the closed neighborhoods of the vertices [11]. They studied several invariants such as the projective dimension, the height, and the Castelnuovo–Mumford regularity and relate them to combinatorial invariants such as the domination number or the matching number. Recently, they continued their study with chordal graphs [7]. Closed neighborhood ideals of graphs have been studied also by Joseph, Roy, and Singh studied the minimal free resolution in the framework of the Barile–Macchia resolutions [6]. They proved that for trees, the Barile–Macchia resolution is minimal. These ideals have been studied also for particular classes of graphs. For instance, Nasernejad, Bandari, and Roberts proved that the closed neighborhood ideals of complete bipartite graphs are normal, therefore they satisfy the (strong) persistence property [9]. Moreover, Chakraborty, Joseph, Roy, and Singh [3] considered the Castelnuovo–Mumford regularity of closed neighborhood ideals of a forest $G$ and obtained that it is equal to the matching number of $G$, fact that was conjectured by Sharifan and Moradi [11, Conjecture 2.11].
In this paper we pay attention to the closed neighborhood ideal of the square of path graph, $NI(P_{n}^{2})$. We determine the height, the bigheight, the projective dimension, and the Castelnuovo–Mumford regularity, and we characterize all closed neighborhood ideals of $P_{n}^{2}$ which are Cohen–Macaulay. Moreover, if we assume that $NI(P_{n}^{2})\subseteq S=k[x_{1},\ldots,x_{n}]$, we prove that $S/NI(P_{n}^{2})$ is sequentially Cohen–Macaulay for $n\geq 7$. We note that Sharifan and Moradi proved a similar result for complete $r$-partite graphs [11, Theorem 2.10 ((vii))]. We also pay attention to the simplicial complex $\mathcal{NI}[P_{n}^{2}]$ whose facet ideal is $NI(P_{n}^{2})$ and prove that it is shellable. Due to its nice structure, we are able to compute its $f$-vector and the reduced Euler characteristic.
@@ 1. Invariants of the closed neighborhood ideals of $P_{n}^{2}$
For a finite, simple graph $G$ with $V(G)=\{x_{1},\ldots,x_{n}\}$ let $S=k[x_{1},\ldots,x_{n}]$ be the polynomial ring over a field $k$. (The same notation will be used for both vertices and variables where there will be no confusion from the context) For a set of vertices $F\subseteq V(G)$ we consider the squarefree monomial $\mathbf{x}_{F}=\prod\limits_{x_{i}\in F}x_{i}$. We will refer to $F$ as the support of the monomial $\mathbf{x}_{F}$ and we will denote it by $\operatorname{supp}(\mathbf{x}_{F})$. Sharifan and Moradi defined the notion of closed neighborhood ideal, denoted by $NI(G)$, as being the squarefree monomial ideal generated by the monomials which correspond to the closed neighborhoods of vertices ([11, Definition 2.1]):
@@
[11, Lemma 2.2] Let $G$ be a graph. Then $\operatorname{ht}(NI(G))=\gamma(G)$ and $\operatorname{bight}(G)=\gamma^{\prime}_{G}$.
@@
[10, Corollary 2.6] If $I$ is a monomial ideal of $S$ and $u$ is a monomial which does not belong to $I$, then the minimal free resolution of $S/(I,u)$ is given by the mapping cone technique provided that there is $x_{i}\in\operatorname{supp}(u)$ such that for all $v\in\mathcal{G}(I)$ $\deg_{x_{i}}(u)>\deg_{x_{i}}(v)$.
@@
We start with the first exact sequence and we determine the Castelnuovo–Mumford regularity and the projective dimension for $\frac{S}{I_{5}(P_{\{2,\ldots,n-1\}})+(x_{1}x_{2}x_{3})}$. By [10, Theorem 2.4] applied in the first exact sequence, we have:
@@ 2. The closed neighborhood ideal of $P_{n}^{2}$ and simplicial complexes
We pay attention now to simplicial complexes in order to determine some other properties of the closed neighborhood ideals. As before, we start bu recalling the notions and results that will be used (for more details. please check [2] or [13].
@@
    If $\Delta$ is a non-pure shellable simplicial complex, then $k[\Delta]$ is sequentially Cohen–Macaulay. (([12],[13, Theorem 6.3.27]))
@@
We recall that a vertex of a simplicial complex $\Delta$ is called a free vertex if it belongs to exactly one facet of $\Delta$. According to [14, Definition 4.1] a simplicial complex $\Delta$ has the free vertex property if it satisfies the following conditions:
@@
[14, Theorem 4.8] Let $\Delta$ be a simplicial complex, $I(\Delta)$ its facet ideal and $SR(\Delta)$ the simplicial complex with the property that $I(\Delta)=I_{SR(\Delta)}$. If $\Delta$ has the free vertex property, then $SR(\Delta)$ is shellable.
@@
[14, Theorem 4.9] If the simplicial complex $\Delta$ on a vertex set $V$ has the free vertex property, then $S/I(\Delta)$ is sequentially Cohen–Macaulay.
@@
    $S/I(P_{n}^{2})$ is sequentially Cohen–Macaulay.
@@
(i) The statement follows easy by Theorem 2.5, Proposition 2.3 and by using the fact that $I(\mathcal{NI}[P_{n}^{2}])=I(P_{n}^{2})$.
@@
[13, Corollary 6.4.20],[8, Corollary 3.33] Let $I$ be a squarefree monomial ideal of $S$ and $S/I$ is sequentially Cohen–Macaulay. Then $\operatorname{pd}(S/I)=\operatorname{bight}(I)$.
We can now determine the big height of the closed neghborhood complex of $P_{n}^{2}$, and the maximal size of a minimal domination set of $P_{n}^{2}$
@@
The first equality holds by Proposition 1.3. The formula follows since $S/NI(P_{n}^{2})$ is sequentially Cohen–Macaulay (Theorem 2.6(i)) and we use Lemma 2.7 and the formulae of projective dimension, Theorem 1.11. ∎
@@
By using [5, Lemma 1.2], we get $I_{\mathcal{NI}[P_{n}^{2}]^{\vee}}=I(\mathcal{NI}[P_{n}^{2}]^{c})=I^{c}(P_{n}^{2})$ and since $\mathcal{NI}[P_{n}^{2}]$ is shellable, the statement follows by Theorem 2.1(ii)(ii). ∎
@@
for all $i=0,\ldots,d$ (here $\dim(\Delta)=d-1$), take into account that in our case $d=5$ and use Proposition 2.10. ∎
@@
For $n\geq 7$ we have $\chi(\mathcal{NI}[P_{n}^{2}])=1$, $\overline{\chi}(\mathcal{NI}[P_{n}^{2}])=0$ and $e(S/\mathcal{NI}[P_{n}^{2}])=n-6$.
@@
Now $\overline{\chi}(\mathcal{N}I[P_{n}^{2}])=\chi(\mathcal{NI}[P_{n}^{2}])-1=0$. For the multiplicity, the formula follows easy by using Corollary 2.11: $e(S/\mathcal{NI}[P_{n}^{2}])=\sum\limits_{i=0}^{d}h_{i}=1+(n-5)-4+2=n-6$. ∎

### v2

# arXiv:2602.14163v2[math.AC] 27 Feb 2026

@@ Abstract.
Keywords: Castelnuovo–Mumford regularity, projective dimension, simplicial complex
MSC: Primary 13D05 Secondary 13F55
@@ Introduction
In 2020, Sharifan and Moradi defined the notion of closed neighborhood ideal of a graph as being the squarefree monomial ideal whose generators are given by the closed neighborhoods of the vertices [12]. They studied several invariants such as the projective dimension, the height, and the Castelnuovo–Mumford regularity and relate them to combinatorial invariants such as the domination number or the matching number. Recently, they continued their study with chordal graphs [8]. Closed neighborhood ideals of graphs have been studied also by Joseph, Roy, and Singh studied the minimal free resolution in the framework of the Barile–Macchia resolutions [6]. They proved that for trees, the Barile–Macchia resolution is minimal. These ideals have been studied also for particular classes of graphs. For instance, Nasernejad, Bandari, and Roberts proved that the closed neighborhood ideals of complete bipartite graphs are normal, therefore they satisfy the (strong) persistence property [10]. Moreover, Chakraborty, Joseph, Roy, and Singh [3] considered the Castelnuovo–Mumford regularity of closed neighborhood ideals of a forest $G$ and obtained that it is equal to the matching number of $G$, fact that was conjectured by Sharifan and Moradi [12, Conjecture 2.11].
In this paper we pay attention to the closed neighborhood ideal of the square of path graph, $NI(P_{n}^{2})$. We determine the height, the bigheight, the projective dimension, and the Castelnuovo–Mumford regularity, and we characterize all closed neighborhood ideals of $P_{n}^{2}$ which are Cohen–Macaulay. Moreover, if we assume that $NI(P_{n}^{2})\subseteq S=k[x_{1},\ldots,x_{n}]$, we prove that $S/NI(P_{n}^{2})$ is sequentially Cohen–Macaulay for $n\geq 7$. A similar result was obtained by Sharifan and Moradi for complete $r$-partite graphs [12, Theorem 2.10 ((vii))]. We also pay attention to the simplicial complex $\mathcal{NI}[P_{n}^{2}]$ whose facet ideal is $NI(P_{n}^{2})$ and prove that it is shellable. Due to its nice structure, we are able to compute its $f$-vector and the reduced Euler characteristic.
@@ 1. Invariants of the closed neighborhood ideals of $P_{n}^{2}$
For a finite, simple graph $G$ with $V(G)=\{x_{1},\ldots,x_{n}\}$ let $S=k[x_{1},\ldots,x_{n}]$ be the polynomial ring over a field $k$. (The same notation will be used for both vertices and variables where there will be no confusion from the context) For a set of vertices $F\subseteq V(G)$ we consider the squarefree monomial $\mathbf{x}_{F}=\prod\limits_{x_{i}\in F}x_{i}$. We will refer to $F$ as the support of the monomial $\mathbf{x}_{F}$ and we will denote it by $\operatorname{supp}(\mathbf{x}_{F})$. Sharifan and Moradi defined the notion of closed neighborhood ideal, denoted by $NI(G)$, as being the squarefree monomial ideal generated by the monomials which correspond to the closed neighborhoods of vertices ([12, Definition 2.1]):
@@
[12, Lemma 2.2] Let $G$ be a graph. Then $\operatorname{ht}(NI(G))=\gamma(G)$ and $\operatorname{bight}(G)=\gamma^{\prime}_{G}$.
@@
[11, Corollary 2.6] If $I$ is a monomial ideal of $S$ and $u$ is a monomial which does not belong to $I$, then the minimal free resolution of $S/(I,u)$ is given by the mapping cone technique provided that there is $x_{i}\in\operatorname{supp}(u)$ such that for all $v\in\mathcal{G}(I)$ $\deg_{x_{i}}(u)>\deg_{x_{i}}(v)$.
@@
We start with the first exact sequence and we determine the Castelnuovo–Mumford regularity and the projective dimension for $\frac{S}{I_{5}(P_{\{2,\ldots,n-1\}})+(x_{1}x_{2}x_{3})}$. By [11, Theorem 2.4] applied in the first exact sequence, we have:
@@ 2. The closed neighborhood ideal of $P_{n}^{2}$ and simplicial complexes
We pay attention now to simplicial complexes in order to determine some other properties of the closed neighborhood ideals. As before, we start by recalling the notions and results that will be used through this section (for more details, please check [2] or [14]).
@@
    If $\Delta$ is a non-pure shellable simplicial complex, then $k[\Delta]$ is sequentially Cohen–Macaulay. (([13], [14, Theorem 6.3.27]))
@@
We recall that a vertex of a simplicial complex $\Delta$ is called a free vertex if it belongs to exactly one facet of $\Delta$. According to [15, Definition 4.1] a simplicial complex $\Delta$ has the free vertex property if it satisfies the following conditions:
@@
[15, Theorem 4.8] Let $\Delta$ be a simplicial complex, $I(\Delta)$ its facet ideal and $SR(\Delta)$ the simplicial complex with the property that $I(\Delta)=I_{SR(\Delta)}$. If $\Delta$ has the free vertex property, then $SR(\Delta)$ is shellable.
The Stanley–Reisner complex of $NI(G)$ is the dominance complex of a graph $G$ which was studied by Matsushita and Wakatsuki [7]. We obtain that, for $P_{n}^{2}$, the dominance complex is sequentially Cohen–Macaulay Theorem 2.6.
@@
[15, Theorem 4.9] If the simplicial complex $\Delta$ on a vertex set $V$ has the free vertex property, then $S/I(\Delta)$ is sequentially Cohen–Macaulay.
@@
    $S/NI(P_{n}^{2})$ is sequentially Cohen–Macaulay.
@@
(i) The statement follows easy by Theorem 2.5, Proposition 2.3 and by using the fact that $I(\mathcal{NI}[P_{n}^{2}])=NI(P_{n}^{2})$.
@@
[14, Corollary 6.4.20],[9, Corollary 3.33] Let $I$ be a squarefree monomial ideal of $S$ and $S/I$ is sequentially Cohen–Macaulay. Then $\operatorname{pd}(S/I)=\operatorname{bight}(I)$.
We can now determine the big height of the closed neighborhood complex of $P_{n}^{2}$, and the maximal size of a minimal domination set of $P_{n}^{2}$
@@
The first equality holds by Proposition 1.3. The formula follows since $S/NI(P_{n}^{2})$ is sequentially Cohen–Macaulay (Theorem 2.6(i)) and we use Lemma 2.7 and the formulae of projective dimension given in Theorem 1.11. ∎
We can also prove that the complementary ideal of $P_{n}^{2}$ has linear quotients.
@@
By using [5, Lemma 1.2], we get $I_{\mathcal{NI}[P_{n}^{2}]^{\vee}}=I(\mathcal{NI}[P_{n}^{2}]^{c})=I^{c}(P_{n}^{2})$ and since $\mathcal{NI}[P_{n}^{2}]$ is shellable, the statement follows by Theorem 2.1(ii). ∎
@@
for all $i=0,\ldots,d$ (here $\dim(\Delta)=d-1$), taking into account that in our case $d=5$ and use Proposition 2.10. ∎
@@
For $n\geq 7$ we have $\chi(\mathcal{NI}[P_{n}^{2}])=1$, $\overline{\chi}(\mathcal{NI}[P_{n}^{2}])=0$ and $e(S/NI(P_{n}^{2}))=n-6$.
@@
Now $\overline{\chi}(\mathcal{N}I[P_{n}^{2}])=\chi(\mathcal{NI}[P_{n}^{2}])-1=0$. For the multiplicity, the formula follows easy by using Corollary 2.11: $e(S/NI(P_{n}^{2})=\sum\limits_{i=0}^{d}h_{i}=1+(n-5)-4+2=n-6$. ∎